OpenStudy (anonymous):

find domain and range of x-2 / x+4

OpenStudy (anonymous):

to find the domain solve the denominator to see where it cannot equal zero, there is a vertical asymptote at that x value thus the domain will skip/not count that x value: \(x+4\neq 0~~\implies~~x\neq-4\) \(domain:~(-\infty,-4)\cup(-4,\infty)\) the range is all real numbers: \((-\infty,\infty)\)

OpenStudy (anonymous):

I thought that too, but the range is 1. I set it equal to 1 and didn't have a solution. I was just looking for another way to solve it

OpenStudy (anonymous):

oh right sorry that is correct...there is a horizontal asymptote at 1 so the range is \((-\infty , 1)\cup(1,\infty)\) ... my bad